# Robust Utility-Preserving Text Anonymization Based on Large Language Models

### v1

@@ Abstract
Text anonymization is crucial for sharing sensitive data while maintaining privacy. Existing techniques face the emerging challenges of re-identification attack ability of Large Language Models (LLMs), which have shown advanced capability in memorizing detailed information and patterns as well as connecting disparate pieces of information. In defending against LLM-based re-identification attacks, anonymization could jeopardize the utility of the resulting anonymized data in downstream tasks—the trade-off between privacy and data utility requires deeper understanding within the context of LLMs. This paper proposes a framework composed of three LLM-based components—a privacy evaluator, a utility evaluator, and an optimization component, which work collaboratively to perform anonymization. To provide a practical model for large-scale and real-time environments, we distill the anonymization capabilities into a lightweight model using Direct Preference Optimization (DPO). Extensive experiments demonstrate that the proposed models outperform baseline models, showing robustness in reducing the risk of re-identification while preserving greater data utility in downstream tasks. [^1] Our code and dataset are available at Github.
@@ 1 Introduction
Privacy protection is a fundamental societal value, enforced through various legal frameworks, e.g., the General Data Protection Regulation (GDPR) in the European Union and the California Consumer Privacy Act (CCPA) in the United States Voigt and Von dem Bussche (2017), among others. The recent advancement in large language models (LLMs) and artificial intelligence (AI) tools present both challenges and opportunities in achieving the goal.
Text anonymization is a critical method for safeguarding private and sensitive information. However, current techniques are vulnerable to disclosure threats from increasingly sophisticated Large Language Models (LLMs). Recent studies have demonstrated that these models can re-identify private information, even from texts anonymized by advanced methods Patsakis and Lykousas (2023); Staab et al. (2024a); Nyffenegger et al. (2024).
@@
The first key challenge and requirement is, therefore, defending against LLM-based re-identification attacks. In defending against these powerful models, the anonymization process may compromise the utility of the resulting anonymized data in downstream tasks Mozes and Kleinberg (2021); Patsakis and Lykousas (2023). As shown in fig. 1, while the current state-of-the-art (SoTA) method, which conducts anonymization based on iterative refining according to feedback from a simulated attacker Staab et al. (2024b), can defend against re-identification attack well, it may eliminate the information crucial for the downstream task. We believe that the trade-off between privacy and data utility requires deeper understanding within the context of LLMs, in which LLMs’ re-identification capacity challenges the existing anonymization models, while if properly utilized, LLMs can help build more capable anonymization components to mitigate the discussed adversaries.
In this paper, we introduce a novel framework named Robust Utility-Preserving Text Anonymization (RUPTA), consisting of a privacy evaluator (P-Evaluator), a utility evaluator (U-Evaluator), and an optimization component.
@@
These components are built on LLMs, where the P-Evaluator assesses re-identification risks and provides guidance to enhance anonymization robustness against re-identification attacks, the U-Evaluator gauges downstream tasks’ performance to indicate the level of preserved utility, and the optimization component iteratively edits the text based on these evaluation results to jointly optimize both objectives until pre-defined conditions are met. As shown in fig. 1, RUPTA can ensure privacy-preserving performance comparable to the SOTA method while retaining critical information necessary for accurately classifying the text pertaining to a Tennis Player.
The anonymization models based on LLMs often rely heavily on time-consuming and resource-intensive interactions with LLMs, making these models less feasible for large-scale or real-time applications. To mitigate this problem, we distill the anonymization capabilities into a lightweight model. Our experiments show that the fine-tuned lightweight model achieves a performance comparable to GPT-4, and utilizing Direct Preference Optimization (DPO) Rafailov et al. (2023) enhances the anonymization efficacy. Our main contributions are summarized as follows:
@@
  • To the best of our knowledge, this is the first work to simultaneously optimize privacy and utility in text anonymization using SoTA LLMs, which is crucial for real-life applications.
@@
  • We propose a novel framework for text anonymization that is built on the powerful ability of LLMs, consisting of a privacy evaluator, utility evaluator, and optimizer component, which work jointly to perform anonymization and show superior performance over the baseline models.
  • We develop more practical methods based on DPO to distill the anonymization capabilities into lightweight models with performance comparable to the teacher models.
@@
  • We create a new dataset using the celebrity biographies from DBpedia Dan (2019) with occupation labels, serving as a practical benchmark for evaluating the impact of anonymization methods on utility. Anonymization results from LLMs are also included to aid future text anonymization research.
@@ 2 Related Work
The task is primarily addressed through natural language processing (NLP) and privacy-preserving data publishing (PPDP) approaches. NLP methods use sequence labeling models trained on manually annotated data to identify and remove pre-defined categories of sensitive entities, such as names and phone numbers Hathurusinghe et al. (2021); Francopoulo and Schaub (2020). Rather than masking entities according to the pre-defined categories, the PPDP-based approaches mask entities according to the disclosure risk calculated through a privacy model defined by domain experts Sánchez and Batet (2016, 2017). However, most existing studies either neglect the utility of anonymized text for downstream tasks or only evaluate it post-anonymization Yermilov et al. (2023); Staab et al. (2024b), complicating the identification of a strategy that optimally balances privacy and utility. Furthermore, commonly used datasets Lebret et al. (2016); Pilán et al. (2022) in this field often lack labels for specific downstream tasks, rendering it difficult to assess the impact of anonymization operations on them.
LLMs as the Black-box Optimizer
Optimization entails the iterative generation and evaluation of solutions to enhance a specific objective function. Leveraging their robust knowledge storage and generation capabilities, LLMs can identify optimal solutions for intricate real-world optimization problems through effective prompting without necessitating additional training Prasad et al. (2023); Zhou et al. (2023). In the context of multi-objective optimization problems (MOPs), which involve two or more conflicting objectives, current methodologies typically combine Evolutionary Algorithms with LLMs Yang and Li (2023). This approach, however, requires numerous objective evaluations, rendering it impractical for scenarios where evaluating objectives is costly. Our proposed RUPTA serves as an alternative when the preference over objectives is pre-defined.
Figure 2: The framework of our proposed RUPTA method
3 Our Approach
@@
In this section, we present our proposed RUPTA framework, which protects the privacy of the sensitive text while maintaining its utility for analytical purposes. The overview of our framework is shown in fig. 2. Given a span of text $\boldsymbol{x}_{0}$, RUPTA iteratively refines the anonymized text to optimize the privacy and utility objectives simultaneously. At iteration $t+1$, the previously anonymized text $\boldsymbol{x}_{t}$ is input into the system, as shown in the bottom left of fig. 2. The privacy evaluator (P-Evaluator) analyzes $\boldsymbol{x}_{t}$ to determine its privacy protection level based on the ground-truth personal information $y$ and provide feedback to enhance its robustness against re-identification attacks. The utility evaluator (U-Evaluator) assesses its usefulness for the downstream tasks based on the corresponding ground-truth label $c$. Feedback from both evaluators is then used by the optimizer to refine the text using available editing operations, producing the updated text $\boldsymbol{x}_{t+1}$, as shown in the top right of fig. 2. Specific content of the involved instructions can be found in section C.1.
@@ 3.1 Problem Formulation
The text anonymization challenge can be recast as a multi-objective optimization problem with two conflicting objectives: privacy and utility. In this context, privacy should be prioritized over utility. This hierarchy is established by ordering the objectives, transforming the problem into a lexicographic optimization issue Zykina (2004). The primary objective is to maximize the level of privacy preservation, ensuring that sensitive information is well-protected against re-identification risks. The secondary objective is to maintain as much useful information as possible in the anonymized text for analytical tasks. This lexicographic optimization problem can be formally expressed as
@@
where $f_{p}(\cdot)$ and $f_{u}(\cdot)$ denote the privacy and utility objective function, respectively. $\mathcal{X}_{0}$ denotes the set of all possible edits of $\boldsymbol{x}_{0}$. A solution $\boldsymbol{x}_{a}\in\mathcal{X}_{0}$ is lexicographically preferable to another solution $\boldsymbol{x}_{b}\in\mathcal{X}_{0}$, denoted as $\boldsymbol{x}_{a}\succ_{\text{lex}}\boldsymbol{x}_{b}$, if and only if $f_{p}(\boldsymbol{x}_{a})>f_{p}(\boldsymbol{x}_{b})$ or $(f_{p}(\boldsymbol{x}_{a})=f_{p}(\boldsymbol{x}_{b})\text{ and }f_{u}(\boldsymbol{x}_{a})>f_{u}(\boldsymbol{x}_{b}))$. To solve this lexicographic optimization problem, we propose RUPTA, an iterative method with LLMs to generate, evaluate, and optimize the anonymized text.
3.2 The Privacy Evaluator
Input Anonymized text $\boldsymbol{y}_{t}$, ground-truth personal information $x$, instruction $\mathbf{I}_{p}$, P-Evaluator $\mathcal{LLM}(\cdot)$
@@ 3.2 The Privacy Evaluator
The role of the Privacy Evaluator (P-Evaluator) is to assess the privacy protection level of the anonymized text, ensuring that private content is adequately obscured against re-identification. Besides, it is essential to provide textual feedback to the LLM optimizer as guidance Pryzant et al. (2023). Thus, the privacy objective evaluation process $f_{p}(\cdot)$ is formally defined as
$\boldsymbol{f}_{t},p_{t}=f_{p}(\boldsymbol{x}_{t})$ (2)
where $p_{t}$ denotes the value of the privacy objective and $\boldsymbol{f}_{t}$ denotes the textual feedback. We describe the detailed process of privacy evaluation in algorithm 1.
P-Evaluator is instantiated as an LLM. Given the anonymized text $\boldsymbol{x}_{t}$, we concatenate it with the privacy inference instruction $\mathbf{I}_{p}$ as input to prompt the P-Evaluator to semantically infer the personal information as shown in line 1 of algorithm 1, where ||||| | denotes concatenation. This step generates top-$K$ inference results $[y^{\prime}_{i}]_{1}^{K}$ for the personal information. Each result is then compared with the ground-truth personal information $y$. If a match is found within these top-$K$ results, its rank is used as the scalar privacy score $p_{t}$. Further, the evaluator is prompted to provide natural language feedback $\boldsymbol{f}_{t}$ detailing the clues that led to the correct inference. Otherwise, we set the $p_{t}$ as $K+1$, representing the maximum achievable score for the privacy objective.
The scalar score $p_{t}$ quantifies the privacy risk associated with the anonymized text, while the textual feedback $\boldsymbol{f}_{t}$ offers qualitative insights, guiding the lexicographic optimizer on how to better obscure identifiable information. The value of $K$ serves as a customizable parameter that adjusts the sensitivity of the privacy evaluation, with higher values indicating a more inclusive search for potential privacy breaches, thus facilitating a manually adjustable trade-off between privacy and utility.
3.3 The Utility Evaluator
The Utility Evaluator (U-Evaluator) is used to ensure that the anonymized text retains its utility for specific analytical tasks, a critical consideration for practical applications across various domains. It analyzes the anonymized text $\boldsymbol{x}_{t}$, specifically assessing its effectiveness in supporting accurate occupation classification $c$. The formal utility objective evaluation process is defined as
@@ 3.3 The Utility Evaluator
In this paper, we instantiate the U-evaluator with an LLM. Given the anonymized text $\boldsymbol{x}_{t}$ and the corresponding ground-truth occupation label $c$, the LLM-based U-evaluator follows the instruction $\mathbf{I}_{u}$ to output a confidence score $u_{t}$
@@
this confidence score quantifies the evaluator’s uncertainty that $\boldsymbol{x}_{t}$ can be correctly classified into the ground truth occupation category $c_{t}$, reflecting the degree to which key utility information is preserved.
To better align feedback with real-world use scenarios, the U-Evaluator can be instantiated with the actual model employed in the downstream task. For example, if the anonymized text is intended for sentiment analysis, the U-Evaluator can be instantiated with a sentiment analysis model. The utility score $u_{t}$ can then be calculated through the logit of the ground-truth label following the traditional uncertainty quantification method Sensoy et al. (2021).
3.4 Lexicographic Optimizer
Lexicographic optimization (LO) is a special case of MOPs where multiple conflicting objectives are to be maximized simultaneously. In LO, objectives are ranked in order of importance, enabling prioritization of the most critical objectives. The LO problem is generally solved by the sequential optimization method Zykina (2004); Zhang et al. (2022). Specifically, regarding the text anonymization problem, privacy and utility are the two objectives, and privacy should be prioritized.
RUPTA employs the LLM as a black-box lexicographic optimizer in a zero-shot manner, where the LLM is prompted to achieve better solutions incrementally based on the history of optimization results and objective evaluations. The overall prompt consists of the pre-defined overall optimization description prompt $\mathbf{I}_{r}$, the memory module $\mathcal{M}$, the meta instruction variable $\boldsymbol{I}_{me}$ and the textual feedback $\boldsymbol{f}_{t}$ from P-Evaluator. The memory module $\mathcal{M}$ stores history optimization results and their corresponding privacy and utility objective values. Formally, we have =$\mathcal{M}=\{(\boldsymbol{x}_{i},p_{i},u_{i},r_{i})|i=1,2,...,t\}$.
To ensure that the primary goal of achieving maximum privacy is prioritized and only after the privacy objective is satisfactorily met does the optimizer focus on improving utility, the lexicographic-optimizer LLM operates in two different modes. When the privacy objective value has not yet reached the pre-set maximum $K+1$, the lexicographic optimizer should focus on maximizing the privacy objective, which is achieved by taking the value of meta instruction variable as $\mathbf{I}_{pr}$ that instructs the LLM to further anonymize $\boldsymbol{x}_{t}$ according to the textual feedback $\boldsymbol{f}_{t}$. The process can be formulated as

### v2

@@ Abstract
Anonymizing text that contains sensitive information is crucial for a wide range of applications. Existing techniques face the emerging challenges of the re-identification ability of large language models (LLMs), which have shown advanced capability in memorizing detailed information and reasoning over dispersed pieces of patterns to draw conclusions. When defending against LLM-based re-identification, anonymization could jeopardize the utility of the resulting anonymized data in downstream tasks. In general, the interaction between anonymization and data utility requires a deeper understanding within the context of LLMs. In this paper, we propose a framework composed of three key LLM-based components: a privacy evaluator, a utility evaluator, and an optimization component, which work collaboratively to perform anonymization. Extensive experiments demonstrate that the proposed model outperforms existing baselines, showing robustness in reducing the risk of re-identification while preserving greater data utility in downstream tasks. We provide detailed studies on these core modules. To consider large-scale and real-time applications, we investigate the distillation of the anonymization capabilities into lightweight models. All of our code and datasets will be made publicly available at Github [^1] https://github.com/UKPLab/acl2025-rupta.
@@ 1 Introduction
Privacy protection is a fundamental societal value, enforced in various legal systems such as the General Data Protection Regulation (GDPR) in the European Union and the California Consumer Privacy Act (CCPA) in the United States Voigt and Von dem Bussche (2017), among many others. The recent advancement in AI and large language models (LLMs) presents both challenges and opportunities for privacy protection.
Anonymization is a critical approach to safeguarding private and sensitive information. However, current techniques are vulnerable to disclosure threats from increasingly sophisticated LLMs. For example, recent studies have demonstrated that such models can re-identify private information, even from texts anonymized by advanced methods Patsakis and Lykousas (2023); Staab et al. (2024a); Nyffenegger et al. (2024).
@@
The first key challenge and requirement is, therefore, defending against LLM-based re-identification attacks. In combating these powerful models, the anonymization process may compromise the utility of the resulting anonymized data in downstream tasks Mozes and Kleinberg (2021); Patsakis and Lykousas (2023). As shown in fig. 1, while the current state-of-the-art (SoTA) method, which conducts anonymization based on iterative refining according to feedback from a simulated attacker Staab et al. (2024b), can defend against re-identification attack well, it may eliminate the information crucial for the downstream task. Existing studies, however, evaluate anonymized text mainly from the perspective of text quality Dou et al. (2023); Staab et al. (2024b), lacking investigation of the impact on downstream tasks. In general, the interactions between remaining privacy and maintaining utility require a deeper understanding within the context of LLMs, where LLMs’ re-identification capacity challenges the safety of existing anonymization models, while if properly utilized, could be leveraged to build more capable anonymization components.
@@
We introduce a framework named Robust Utility-Preserving Text Anonymization (RUPTA), consisting of a privacy evaluator (P-Evaluator), a utility evaluator (U-Evaluator), and an optimization component. These components are built on LLMs, where the P-Evaluator assesses re-identification risks and provides guidance to enhance anonymization robustness, the U-Evaluator gauges downstream tasks’ performance to indicate the level of preserved utility, and the optimization component iteratively edits the text based on these evaluations to optimize both objectives until the pre-defined conditions are met. RUPTA outperforms existing baselines based on LLMs, showing robustness in reducing the risk of re-identification while preserving greater data utility in downstream tasks. Note that the privacy protection level can be customized in the proposed framework. Since the anonymization based on LLMs could be time-consuming and resource-intensive, we additionally investigated the distillation of the anonymization capabilities into lightweight models. Our main contributions are summarized as follows:
@@
  • To the best of our knowledge, this is the first work to provide comprehensive studies on anonymization and utility in the setup of LLMs, which are crucial for real-world applications.
@@
  • We propose a novel framework for text anonymization that is built on the powerful ability of LLMs, consisting of a privacy evaluator, a utility evaluator, and an optimizer component. They work in tandem and show superior performance over the existing models. We provide detailed studies on these core modules. To provide a practical model for real-time environments, we investigate the distillation of anonymization capabilities into smaller models.
@@
  • We create a new dataset using the celebrity biographies from DBpedia Dan (2019) with occupation labels, serving as a practical benchmark for evaluating the impact of anonymization methods on downstream tasks. Anonymization results from LLMs are also included to aid future text anonymization research.
@@ 2 Related Work
To keep privacy when sharing sensitive data, text anonymization serves as a critical alternative to differential privacy-based methods Feyisetan et al. (2019); Xu et al. (2020); Mattern et al. (2022) and representation learning-based methods Coavoux et al. (2018) for its high fidelity. This task is primarily addressed through techniques from natural language processing (NLP) and privacy-preserving data publishing (PPDP) Lison et al. (2021). NLP techniques generally employ sequence labeling models, which are trained on manually annotated datasets to identify and obscure predefined categories of sensitive entities such as names and phone numbers Hathurusinghe et al. (2021); Francopoulo and Schaub (2020). In contrast, PPDP methods obscure entities based on a disclosure risk assessed through a privacy model, which is defined by domain experts—examples include C-sanitize Sánchez and Batet (2016, 2017). However, most existing studies neglect the utility of anonymized text for downstream tasks or perform post-anonymization evaluations focused on text quality Yermilov et al. (2023); Staab et al. (2024b), compromising the flexibility of strategies that are able to consider both privacy and utility.
Furthermore, commonly used datasets Lebret et al. (2016); Pilán et al. (2022) often lack labels for specific downstream tasks, making it difficult to assess the impact of anonymization operations on them.
Anonymization in the Context of LLMs.
Significant advancements in LLMs have introduced both challenges and opportunities for text anonymization. Prior to the advent of LLMs, many studies on text anonymization Sánchez and Batet (2016, 2017) focused on potential re-identification risks posed by adversaries that could exploit extensive external background knowledge, such as information available on the Web. These studies typically relied on relatively simple re-identification methods that have low attack success rates, such as analyzing the (co-)occurrence counts of terms on the web or examining lexical and taxonomic relationships. With the rapid development of LLM’s reasoning abilities, Staab et al. (2024a) and Patsakis and Lykousas (2023) for the first time demonstrated that LLMs can re-identify anonymized text with remarkable accuracy and speed.  Staab et al. (2024b) attempted to harness the capabilities of LLMs to defend against re-identification attacks facilitated by LLMs themselves through deploying an simulated adversarial LLM and using its feedback to inform the anonymization process. Based on this work, Frikha et al. (2024) conducted anonymization by introducing synthetic information to mislead the attacker and introduced an early stopping mechanism to mitigate the deterioration of the utility of the anonymized text. Additionally, Dou et al. (2023) explored interactive anonymization methods that involve fine-tuning LLMs. While these approaches have demonstrated promising performance, their impact on the utility of anonymized text for downstream tasks remains underexplored.
Figure 2: An overview of the proposed RUPTA framework. $\mathbf{x}_{t}$ and $\mathbf{x}_{t+1}$ denote the input and output text in one iteration; $y$ denotes the ground-truth personal information; and $\boldsymbol{f}_{t}$, $[y_{i}^{\prime}]_{1}^{K}$ and $p_{t}$ are the inference feedback, inferred personal information from P-Evaluator and the value of the privacy objective. The ground-truth downstream task label is denoted as $c$, while $u_{t}$ is the value of the utility objective. $\mathcal{M}$ denotes the history optimization results. $\mathbf{I}_{u},\mathbf{I}_{p},\mathbf{I}_{r},\mathbf{I}_{me}$ and $\mathbf{I}_{pa}$ are the prompts used for each component of the method.
3 Our Methods
We present the Robust Utility-Preserving Text Anonymization (RUPTA) framework, aiming to protect the privacy of sensitive text against the re-identification attack from LLMs while maintaining its utility for downstream tasks.
@@ 3 Our Methods
The overview of RUPTA is depicted in fig. 2. Given a span of text $\boldsymbol{x}_{0}$, RUPTA iteratively refines the text to optimize the privacy and utility objectives simultaneously. At iteration $t+1$, the previously anonymized text $\boldsymbol{x}_{t}$ is taken as input, as shown in the bottom left of the figure. The privacy evaluator (P-Evaluator) analyzes $\boldsymbol{x}_{t}$ to determine its privacy protection level based on the ground-truth personal information $y$ and then provides feedback to enhance its robustness against re-identification attacks. The utility evaluator (U-Evaluator) assesses its usefulness for the downstream tasks based on the corresponding ground-truth label $c$. As shown in the top-right part of the figure, feedbacks from both evaluators are consequently used by the optimizer to refine the text using available editing operations, producing the updated text $\boldsymbol{x}_{t+1}$. Details of the involved instructions can be found in section G.2.
@@ 3.1 Problem Formulation
We formulate anonymization as a multi-objective optimization problem, taking into account two objectives: privacy protection and utility of anonymized text. Specifically, this is formulated as Lexicographic Optimization (LO) task Zykina (2004) in which we order the two objectives by giving privacy a higher priority—the primary objective is to maximize the level of privacy protection, ensuring that sensitive information is well-protected against re-identification risks. The secondary objective is to preserve as much useful information as possible in the anonymized text for analytical tasks. The optimization problem can be formally expressed as follows:
@@
where $f_{p}(\cdot)$ and $f_{u}(\cdot)$ denote the privacy and utility objective function, respectively. $\mathcal{X}_{0}$ denotes the set of all possible edits of $\boldsymbol{x}_{0}$. A solution $\boldsymbol{x}_{a}\in\mathcal{X}_{0}$ is lexicographically preferable to another solution $\boldsymbol{x}_{b}\in\mathcal{X}_{0}$, denoted as $\boldsymbol{x}_{a}\succ_{\text{lex}}\boldsymbol{x}_{b}$, if and only if $f_{p}(\boldsymbol{x}_{a})>f_{p}(\boldsymbol{x}_{b})$ or $(f_{p}(\boldsymbol{x}_{a})=f_{p}(\boldsymbol{x}_{b})\text{ and }f_{u}(\boldsymbol{x}_{a})>f_{u}(\boldsymbol{x}_{b}))$. To solve this lexicographic optimization problem, RUPTA takes an iterative method based on LLMs to generate, evaluate, and optimize the anonymized text. The details of our LO module are discussed below in Section 3.4.
3.2 The P-Evaluator
The role of our Privacy Evaluator (P-Evaluator) is to assess the privacy protection level of the anonymized text, ensuring that private content is adequately obscured against re-identification. It is essential to provide textual feedback to the LLM optimizer as guidance Pryzant et al. (2023). Thus, the privacy objective evaluation process $f_{p}(\cdot)$ is formally defined as
$\boldsymbol{f}_{t},p_{t}=f_{p}(\boldsymbol{x}_{t})$ (2)
where $p_{t}$ denotes the value of the privacy objective and $\boldsymbol{f}_{t}$ denotes the textual feedback. We depict the detailed process of P-Evaluation in algorithm 1.
Input Anonymized text $\boldsymbol{x}_{t}$, ground-truth personal information $y$, instruction $\mathbf{I}_{p}$, P-Evaluator $\mathcal{LLM}(\cdot)$
@@ 3.2 The P-Evaluator
Algorithm 1 Privacy Objective Evaluation $f_{p}$
P-Evaluator is instantiated as an LLM. Given the anonymized text $\boldsymbol{x}_{t}$, we concatenate it with the privacy inference instruction $\mathbf{I}_{p}$ as input to prompt the P-Evaluator to semantically infer the personal information as shown in line 1 of algorithm 1, where ||||| | denotes concatenation. This step generates top-$K$ re-identification results $[y^{\prime}_{i}]_{1}^{K}$ for the personal information. Each result is then compared with the ground-truth personal information $y$. If a match is found within these top-$K$ results, its rank is used as the scalar privacy score $p_{t}$. Further, the evaluator is prompted to provide natural language feedback $\boldsymbol{f}_{t}$, detailing the clues that led to the correct inference. Otherwise, we set $p_{t}$ as $K+1$, representing the maximum score for the privacy objective.
The score $p_{t}$ quantifies the privacy risk associated with the anonymized text, while the textual feedback $\boldsymbol{f}_{t}$ offers qualitative insights, guiding the optimizer to better obscure identifiable information. Note that the value of $K$ serves as a parameter that adjusts the sensitivity of the privacy evaluation, with higher values indicating a more inclusive search for potential privacy breaches, thus facilitating a customizable privacy protection level.
3.3 The U-Evaluator
The Utility Evaluator (U-Evaluator) is designed to ensure that the anonymized text retains its utility for downstream analytical tasks, a crucial consideration for practical applications. It analyzes the anonymized text $\boldsymbol{x}_{t}$, assessing its effectiveness in supporting the ground-truth label $c$. The formal utility objective evaluation process is defined as
@@ 3.3 The U-Evaluator
We instantiate the U-evaluator with an LLM. Given the anonymized text $\boldsymbol{x}_{t}$ and the corresponding ground-truth label $c$, the LLM-based U-evaluator follows the instruction $\mathbf{I}_{u}$ to output a confidence score $u_{t}$:
@@
which quantifies the evaluator’s uncertainty about whether $\boldsymbol{x}_{t}$ can be correctly related into the ground truth label $c$, reflecting the degree to which key utility information is preserved. Note that RUPTA is flexible in that the U-Evaluator can be instantiated with the actual model employed in the downstream task. For example, in applications where anonymized text is intended to be used for sentiment analysis (SA), the U-Evaluator can be instantiated with an SA model. The utility score $u_{t}$ can be calculated using the logit of the ground-truth label following traditional uncertainty quantification methods Sensoy et al. (2021).
3.4 The Optimizer
Lexicographic optimization (LO) is a special case of multi-objective optimization problems where multiple conflicting objectives are to be optimized simultaneously. Again, the general objective of LO has been given above in Section 3.1, where the sub-objectives are ranked in order of importance, enabling the prioritization of the more critical objectives. The LO solver is often built on sequential optimization methods Zykina (2004); Zhang et al. (2022). In our text anonymization problem, we prioritize privacy over utility.
RUPTA employs an LLM as the lexicographic optimizer by prompting it iteratively to acquire solutions based on the history of optimization results and objective evaluation results. The overall prompt consists of the pre-defined optimization description prompt $\mathbf{I}_{r}$, the memory module $\mathcal{M}$, the meta instruction variable $\boldsymbol{I}_{me}$ and the textual feedback $\boldsymbol{f}_{t}$ from P-Evaluator. The memory module =$\mathcal{M}=\{(\boldsymbol{x}_{i},p_{i},u_{i})|i=1,2,...,t\}$ stores history optimization results and their corresponding privacy and utility objective values.
To ensure that the primary objective of achieving maximum privacy is prioritized over utility, the lexicographic-optimizer LLM operates in two different modes. When the privacy objective value has not yet reached the pre-set maximum $K+1$, the lexicographic optimizer should focus on maximizing the privacy objective, which is achieved by taking the value of the meta-instruction variable as $\mathbf{I}_{pr}$ that instructs the LLM to further anonymize $\boldsymbol{x}_{t}$ according to textual feedback $\boldsymbol{f}_{t}$. The process can be formulated as